FInd the cube roots of -8i (Answers may be left in trigonometric form). Having trouble solving this
Answer by stanbon(75887)

FInd the cube roots of -8i (Answers may be left in trigonometric form).
-----------------
0-8i
-----
r = sqrt(0^2+8^2) = 8
---
theta = arctan(-8/0) = (3/2)pi
-----------------------------------
(-8i)^(1/3) = r(cis((3/2)pi + npi)/3
-----
Let n = 0 : (-8i)^(1/3) = 8cis(pi/2)
Let n = 1 : (-8i)^(1/3) = 8cis((3/2)+(1/3))pi) = 8cis((11/6)pi)
Let n = 2 : (-8i)^(1/3) = 8cis[(3/2) + (2/3)]pi= 8cis[(13/6)pi]
================
Cheers,
Stan H.
-----------------